If the average service time of a single-server system is 10 minutes. What is the service rate per shift, if each shift is 8 hours?
Answer #1

Average service time = 10 minutes

Each shift is 8 hours = (8×60) minutes = 480 minutes

So service rate = 480/10 = 48 per shift